OpenStudy (anonymous):

Evaluate: -32^0.8 what would the 0.8 be in a radical?

OpenStudy (anonymous):

hmm i forgot but i think so you might wanna check but i think so fractional exponents would also be in a radical form also usually used in calc

OpenStudy (ash2326):

\[ -32^{0.8}\] 0.8=4/5 \[-32^{4/5}\] or \[(-32^4)^{1/5}\] in radical from \[\sqrt[5]{-32^4}\]

OpenStudy (ash2326):

now 32=2^5 so \[(-32)^{1/5}=-2\] we'll have now \[-2^4\] which is -16

OpenStudy (ash2326):

sorry it'd be +16